OpenStudy (spring98):

Medals!! Simplify 12^3/12^7.

OpenStudy (javalos01):

ik

OpenStudy (javalos01):

it

OpenStudy (spring98):

?

rishavraj (rishavraj):

\[\frac{ 1 }{ a^2 } = a^{-2}~~and~~ a^x \times a^y = a^{x+ y}\]

OpenStudy (spring98):

so how do i do it?

rishavraj (rishavraj):

\[\frac{ 1 }{ 12^{7} } = 12^{-7}\]

OpenStudy (cggurumanjunath):

a^x / a^y =a^(x-y)

OpenStudy (jhannybean):

\[\frac{12^3}{12^7} = 12^{3-7} = 12^?\]

OpenStudy (cggurumanjunath):

a=12,x=3,y=7

OpenStudy (jhannybean):

\[\large 12^{-?} \iff \frac{1}{12^?}\]

OpenStudy (spring98):

4

rishavraj (rishavraj):

so it is \[\frac{ 1 }{ 12^4 } \] now wht??

rishavraj (rishavraj):

remember \[\frac{ 1 }{ a } = a^{-1}\]

OpenStudy (spring98):

oh so would it be - 1/12^4?

OpenStudy (anonymous):

12^4

OpenStudy (anonymous):

that would be you answer

rishavraj (rishavraj):

nah @yinkim52001

rishavraj (rishavraj):

@Spring98 u done???

OpenStudy (anonymous):

yah why not?

OpenStudy (anonymous):

there's no negative exponents .

rishavraj (rishavraj):

@yinkim52001 remember \[\frac{ 1 }{ a } = a^{-1}\]

OpenStudy (jhannybean):

You never leave your answers with negative exponents unless it's called for. Therefore \[\frac{1}{12^4} = \frac{1}{12\cdot 12\cdot 12\cdot 12}\] Whatever that is is your answer.

OpenStudy (spring98):

12^4 is the answer right?

OpenStudy (jhannybean):

No.

OpenStudy (anonymous):

no its 1/12^4

OpenStudy (anonymous):

i see why its 1/12^4

rishavraj (rishavraj):

@Jhannybean its asking to just simplfy it ..... so i think we may just write in exponential form

OpenStudy (jhannybean):

You can have various simplifications @rishavraj either in positive exponential form, fully expanded form, or even negative exponent form. They would all be correct.

OpenStudy (spring98):

what is the answer then ? 1/12^4?

OpenStudy (anonymous):

\[\frac{ 1 }{ 12^{4} }\] thats your answer spring

OpenStudy (anonymous):

yup

OpenStudy (jhannybean):

\[\frac{12^{-3}}{12^7} = 12^{-4} =\frac{1}{12^4} = \frac{1}{20736}\] TRy solving the next one yourself @Spring98

OpenStudy (spring98):

thanks @yinkim52001

OpenStudy (anonymous):

No problem